OpenStudy (anonymous):

What is the slope of the line passing through (-3, 5) and (5, -3)? slope = -1 slope = 0 slope = 1 no slope

OpenStudy (eyust707):

Slope is the change in y over the change in x

OpenStudy (4n1m0s1ty):

Hint: Use this eq. (y2 - y1) = k(x2-x1)

OpenStudy (4n1m0s1ty):

where k is slope

OpenStudy (anonymous):

the second one rite

OpenStudy (eyust707):

nope... slope = (y2-y1)/(x2-x1)

OpenStudy (4n1m0s1ty):

So lets see: Point 1: (-3, 5) Point 2: (5, -3) (-3 - 5) = k(5 - -3) k = (-3 - 5)/(5 - -3)

OpenStudy (4n1m0s1ty):

Its basically the change in Y over the change in X. Your change in Y is the difference between the second and first Y coordinates( thus Y2 -Y1), and the change in X is the difference between the second and first X coordinates (X2 - X1).

OpenStudy (anonymous):

-1

OpenStudy (4n1m0s1ty):

Yep, thats it